# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 7.3. Volume preserving map
This subsection aims at proving that there is a volume preserving map between dual hives $H(\alpha,\beta,\gamma,\infty)$ and hives $P_{\alpha,\beta,\gamma}^{g}$ of Definition 2.3.
@@
(7.12) $S^{C}=\{v_{4}:l=(v_{1},v_{2},v_{3},v_{4})\text{ is a }m\text{ lozenge}\}\subset R_{d,n}$
@@
The choice of $v_{4}$ and the coloring on the boundary ensures that $v_{4}$ is never on the boundary of $R_{d,n}$. We will assign to any $(C,L)\in\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$ an element in $\tilde{P}^{C}_{\alpha,\beta,\gamma}$ :
@@
where $\Psi^{C}[L](v)=0$ on the south-east corner of $R_{d,n}$ and adding $L(e)$ when crossing edges of type $1,2$ and $1-L(e)$ for edges of type 00 as in Figure
Figure 32. Values at vertices when traversing an edge $e=(u,v)$.
That the map $\Psi^{C}$ is well defined is due to the face summation constraint $L(e_{1})+L(e_{2})=1-L(e_{0})$ around every face $f\in F_{n}$ having edges $(e_{0},e_{1},e_{2})$ of respective types $0,1,2$ for the label maps $L$ of dual hives in $\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$.
Remark 7.8 (Extension of function of $\tilde{P}^{C}_{\alpha,\beta,\gamma}$ ).
@@
Proposition 7.9 (From color maps to regular labelings).
@@
where the value at vertex $v\in R_{d,n}$ is set as follows. If $v=A_{0}$ the south-east most vertex in $R_{n,d}$, set $g[C](v)=0$. Orient the edges in $E_{d,n}$ around north-pointing faces clockwise and edges around south-pointing faces counterclockwise. For any oriented edge $e=(u,v)$, set
@@
Then, $C\mapsto g[C]$ is a bijection between color maps on $E_{n,d}$ and regular labelings on $R_{n,d}$. For any $g$ regular labeling, its inverse is given by
$C[g]:E_{d,n}$$\rightarrow\{0,1,3,m\}$
@@
where
(7.17) $C[g](e)=\begin{cases}1&\text{ if }g(v)-g(u)=1\\ 0&\text{ if }g(v)-g(u)=2\\ 3&\text{ if }g(v)=g(u)\text{ and }F(e)\text{ has cyclic difference }(1,2,0)\\ m&\text{ if }g(v)=g(u)\text{ and }F(e)\text{ has cyclic difference }(2,1,0).\end{cases}$
@@
Let us show that $g[C]$ is well defined. Since $C$ is a color map, the only colors around a triangular face in $E_{n,d}$ are up to cyclic permutations $(0,0,0),(1,1,1),(1,0,3)$ and $(0,1,m)$. One checks that summing the clockwise differences of values of $g$ going from a vertex to itself around any such color triple gives a zero contribution in $\mathbb{Z}_{3}$. Therefore, the value of $g[C](v)$ does not depend on the choice of the edge $e=(u,v)\in E_{n,d}$ since any different neighbor choice $u^{\prime}$ such that $(u^{\prime},v)\in E_{n,d}$ differ by two triangular faces, see Figure 33 where $(X,Y,Z)$ denote the differences of values of $g$. We have $g(u)+X=g(u^{\prime})+Y^{\prime}[3]$ since $g(u)+X=g(u)-Y-Z=g(w)-Z=g(w)+X^{\prime}+Y^{\prime}=g(u^{\prime})+Y^{\prime}$.
Figure 33. The equality $g(u)+X=g(u^{\prime})+Y^{\prime}[3]$
That $(g[C]^{A},g[C]^{B},g[C]^{C})$ has the right boundary conditions is due to the fact that $C$ is regular. It remains to check the lozenge condition on $g[C]$ from Definition 2.1. Take any lozenge $l=(v^{1},v^{2},v^{3},v^{4})$ and suppose that $g[C](v^{2})=g[C](v^{4})$. Note that from Figure 3, the edge between $v^{2}$ and $v^{4}$ is always oriented from $v^{4}$ to $v^{2}$. The edge $e=(v^{4},v^{2})$ has color either $3$ or $m$. Since $C$ is a color map, the two faces adjacent to $e$ have either $(1,0,3)$ or $(0,1,m)$ colors. The face with vertices $(v^{1},v^{2},v^{4})$, respectively $(v^{3},v^{2},v^{4})$ is always north pointing, respectively south pointing, see Figure 3. If $C(e)=3$, $g[C](v^{1})=g[C](v^{2})+1$ and $g[C](v^{3})=g[C](v^{2})+2$ whereas if $C(e)=m$, $g[C](v^{1})=g[C](v^{2})+2$ and $g[C](v^{3})=g[C](v^{2})+1$. In both cases, $\{g[C](v^{1}),g[C](v^{3})\}=\{g[C](v^{2})+1,g[C](v^{2})+2\}$ and thus $g[C]$ is a regular labeling.
Let us show that $g\mapsto C[g]$ maps a regular labeling $g$ to a color map. Since $g$ is regular, $C[g]$ also is by the same argument as above. Let us show that the only cyclic colors configurations around any triangular face are (0,0,0),(1,1,1,),(1,0,3)(0,0,0),(1,1,1,),(1,0,3)( 0 , 0 , 0 ) , ( 1 , 1 , 1 , ) , ( 1 , 0 , 3 ) and $(0,1,m)$. Take any triangular face and denote $X,Y,Z$ the clockwise differences of values of $g$. Then, $X+Y+Z=0[3]$ so that $(X,Y,Z)\in\{(1,1,1),(2,2,2),(0,1,2),(0,2,1)\}$. Note that we exclude $(0,0,0)$ by the lozenge condition in Definition 2.1 since a rigid lozenge cannot have three vertices with equal values. These possible height differences give the clockwise colors $\{(1,1,1),(0,0,0),(3,1,0),(m,0,1)\}$ respectively. Therefore, $C[g]$ is a color map and by construction $g\mapsto C[g]$ is the inverse of $C\mapsto g[C]$. ∎
@@
The image $\Psi^{C}[(C,L)]$ of any limit hive $(C,L)\in H^{C}(\alpha,\beta,\gamma,\infty)$ can be extended by Remark 7.8 to a function $f:R_{d,n}\rightarrow\mathbb{R}$ such that by construction $\Psi^{C}[(C,L)]=f_{|R_{d,n}\setminus S^{C}}=f_{|Supp(g[C])}$. Let us check that $f\in P^{g[C]}_{\alpha,\beta,\gamma}$. By definition of the extension, $f$ satisfies the equality constraints over any rigid lozenge in $R_{n,d}$. For any other lozenge $l=(v_{1},\dots,v_{4})$, the inequality $f(v_{2})+f(v_{4})\geq f(v_{1})+f(v_{3})$ is equivalent to the inequality (2c) of Definition 7.1.
@@
For $C=C_{0}$, enumerate $e_{1},\ldots,e_{D}$ the horizontal edges in $C_{0}$ (except for the ones on the east band).
$[\Psi^{C_{0}}\circ\Phi^{C_{0}}(u)](v)=\sum_{e_{i}}(1-u_{i})$
@@
If $C$ is general, introduce for each hexagon rotation $C\rightarrow C^{\prime}$ the map $\tilde{R}_{C\rightarrow C^{\prime}}:\tilde{P}^{C}_{\alpha,\beta,\gamma}\rightarrow\tilde{P}^{C^{\prime}}_{\alpha,\beta,\gamma}$ by
@@
    Let us describe how the hexagon rotation $C\rightarrow C^{\prime}$ maps $f$ to another function $f^{\prime}:R_{n,d}\rightarrow\mathbb{R}$. The value of the center vertex $c$ of $h$ is uniquely determined by the position of the rigid lozenge in $h$ and the values of $f$ on $\partial h$. Indeed, if $(v,v^{\prime},v^{\prime\prime})$ are the three other vertices of the rigid lozenge such that $C((c,v^{\prime\prime}))=m$, then $f(c)=f(v)+f(v^{\prime})-f(v^{\prime\prime})$. Note that $(v,v^{\prime},v^{\prime\prime})\in\partial h^{3}$. For every vertex $u\in R_{n,d}$ other that the center vertex $c$ of $h$, we set $f^{\prime}(u)=f(u)$. In the hexagon rotation $C\rightarrow C^{\prime}$, the position of the the rigid lozenge changes and we set $f^{\prime}(c)=f(w)+f(w^{\prime})-f(w^{\prime\prime})$ where $(v,v^{\prime},v^{\prime\prime})\in\partial h^{3}$ are the new vertices of the rigid lozenge on $\partial h^{\prime}=\partial h$.
@@
    The map $\tilde{R}_{C\rightarrow C^{\prime}}(f)$ is defined as the restriction of $f^{\prime}$ of the previous step to $R_{n,d}\setminus S^{C^{\prime}}$.
Note that the map $\tilde{R}_{C\rightarrow C^{\prime}}$ is an affine bijection with integer coefficients whose inverse is given by $\tilde{R}_{C^{\prime}\rightarrow C^{\prime}}$. Let us check that the following diagram is commutative
@@
Let $H=(C,L)\in\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$ having an ABC hexagon $h$ with center vertex $c$. Denote $C^{\prime}$ the color map obtained after any rotation $h\mapsto h^{\prime}$. For every vertex $u\neq c\in R_{n,d}$, one has $\tilde{R}_{C\rightarrow C^{\prime}}(\Psi^{C})(u)=\Psi^{C}(u)$. Moreover, if $u\neq c$ then one can find a path of edges from the south-east most vertex of $R_{n,d}$ to $u$ without any edge incident to $c$. Since the labels of these edges are not changed by $Rot[C\rightarrow C^{\prime}]$, $\Psi^{C^{\prime}}(L^{\prime})(u)=\Psi^{C}(L)(u)=\tilde{R}_{C\rightarrow C^{\prime}}(\Psi^{C}(L))(u)$ as desired. It remains to check that the same property holds for $u=c$. Denote $(v,v^{\prime},v^{\prime\prime})$, respectively $(w,w^{\prime},w^{\prime\prime})$ the vertices on $\partial h$ such that $(c,v,v^{\prime},v^{\prime\prime})$, respectively $(w,w^{\prime},w^{\prime\prime})$, are vertices of the rigid lozenge in $h$, respectively $h^{\prime}$ and that $C((u,v^{\prime\prime}))=C^{\prime}((u,w^{\prime\prime}))=m$. Then,
@@
Since the values of $\Psi$ do not depend on the chosen path, let us choose the following four paths. Take any path $p=(e_{1},\dots,e_{r})\in(E_{n,d})^{r}$ from the south-east origin to $w$ such that for each $1\leq i\leq r$, $e_{i}$ is not an interior edge of $h$. then,
@@
    To evaluate $\Psi^{C^{\prime}}(L^{\prime})(u)$, we append $(w,u)$ to $p$
@@
$\Psi^{C}(L)(w)$ $=\sum_{1\leq i\leq r}L(e_{i})$
$\Psi^{C}(L)(w^{\prime\prime})$ $=\sum_{1\leq i\leq r}L(e_{i})+L((w,w^{\prime\prime}))$
$\Psi^{C}(L)(w^{\prime})$ $=\sum_{1\leq i\leq r}L(e_{i})+L((w,w^{\prime\prime}))+L((w^{\prime\prime},w^{\prime}))$
$\Psi^{C^{\prime}}(L^{\prime})(u)$ $=\sum_{1\leq i\leq r}L^{\prime}(e_{i})+L^{\prime}((w,u)).$
@@
$L^{\prime}((w,u))=L((w,w^{\prime\prime}))+L((w^{\prime\prime},w^{\prime}))-L((w,w^{\prime\prime}))=L((w^{\prime\prime},w^{\prime}))$
@@
$L^{\prime}((w,u))=L((w,w^{\prime\prime}))$
Notice that $(w,u),(w^{\prime\prime},w^{\prime})$ are two edges of the same type in the rigid lozenge in $C^{\prime}$ which implies that $L^{\prime}((w,u))=L^{\prime}((w^{\prime},w^{\prime}))=L((w^{\prime},w^{\prime\prime}))$ where the last equality is due to the fact that $(w^{\prime},w^{\prime\prime})\in\partial h$ so its label is unchanged by $Rot[C,C^{\prime}]$. The commutativity of (7.22) is thus showed.
@@
By Theorem 3.8 we had
(7.23) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=\frac{\Delta(\mathrm{e}^{2i\pi\gamma})}{\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\lim_{N\rightarrow\infty}N^{-(n-1)(n-2)/2}c_{\lambda_{N},\mu_{N}}^{\nu_{N},d}.$
@@
(7.24) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ $=\frac{\Delta(\mathrm{e}^{2i\pi\gamma})}{\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{C}Vol[u\in\mathbb{R}^{D},\Phi^{C}[u]\in H^{C}(\alpha,\beta,\gamma,\infty)]$
(7.25) $=\frac{\Delta(\mathrm{e}^{2i\pi\gamma})}{\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{C}Vol(P^{g[C]}_{\alpha,\beta,\gamma})$
(7.26) $=\frac{\Delta(\mathrm{e}^{2i\pi\gamma})}{\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{g:R_{d,n}\rightarrow\mathbb{Z}_{3}\text{ regular}}Vol_{g}(P^{g}_{\alpha,\beta,\gamma}).$
@@
From [31, Eq. (4.116)], we have

### v2

@@ 7.3. Volume preserving map
This subsection aims at proving that there is a volume preserving map between dual hives $H(\alpha,\beta,\gamma,\infty)$ and hives $P_{\alpha,\beta,\gamma}^{g}$ of Definition 2.4. Refer to the notations of Section 2 for the hives and related notions.
Let $H=(C,L)\in\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$. We assign to $H$ an function $\Psi^{C}:R_{d,n}\rightarrow\mathbb{R}$ constructed as follows. Set $\Psi^{C}[L](v)=d$ where $v$ is the south-east vertex of $R_{d,n}$. For any other vertex $v\in R_{d,n}$ such that $e=(u,v)\in E_{n,d}$ and $\Psi^{C}(u)$ has been set, the value $\Psi^{C}(v)$ is given by
(7.12) $\Psi^{C}(v)=\begin{cases}\Psi^{C}(u)+L(e)&\text{ if $e$ is of type $1$ or $2$ }\\ \Psi^{C}(u)+1-L(e)&\text{ if $e$ is of type $0$. }\end{cases}$
See Figure 28 for a picture of the recursive construction of $\Phi^{C}$ along edges.
Figure 28. Values at vertices when traversing an edge $e=(u,v)$.
@@
(7.13) $S^{C}=\{v_{4}:l=(v_{1},v_{2},v_{3},v_{4})\text{ is a rigid lozenge}\}\subset R_{d,n}$
@@
Moreover, define
@@
where $\Psi^{C}$ is given by (7.12) in the above construction.
Remark that the choice of $v_{4}$ and the coloring on the boundary ensures that $v_{4}$ is never on the boundary of $R_{d,n}$ in the above definition. That the map $\Psi^{C}$ is well defined is due to the face summation constraint $L(e_{1})+L(e_{2})=1-L(e_{0})$ around every face $f\in F_{k}$ having edges $(e_{0},e_{1},e_{2})$ of respective types $0,1,2$ for the label maps $L$ of dual hives in $\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$. Remark that $\Psi^{C}$ depends on $C$ only through its domain and target space.
Remark 7.8 (Extension of functions of $\tilde{P}^{C}_{\alpha,\beta,\gamma}$ ).
@@
where the value at vertex $v\in R_{d,n}$ is set as follows. If $v=A_{0}$ the south-east most vertex in $R_{d,n}$, set $g[C](v)=0$. Orient the edges in $E_{n,d}$ around direct triangles clockwise and edges around reversed triangles counterclockwise. For any oriented edge $e=(u,v)$, set
@@
Proposition 7.9 (From color maps to regular labelings).
The above map $C\mapsto g[C]$ is a bijection between color maps on $E_{n,d}$ and regular labelings on $R_{d,n}$. For any regular labeling $g$, its inverse is given by
$C[g]:E_{n,d}$$\rightarrow\{0,1,3,m\}$
@@
where, if $w\in R_{d,n}$ denotes the third vertex so that $(u,v,w)$ is a direct triangular face,
(7.18) $C[g](e)=\begin{cases}1&\text{ if }g(v)-g(u)=1\\ 0&\text{ if }g(v)-g(u)=2\\ 3&\text{ if }g(v)=g(u)\text{ and }g(w)=g(u)-1=g(v)-1\\ m&\text{ if }g(v)=g(u)\text{ and }g(w)=g(u)+1=g(v)+1.\end{cases}$
@@
Let us show that $g[C]$ is well defined. Since $C$ is a color map, the only colors around a triangular face in $E_{n,d}$ are up to cyclic permutations $(0,0,0),(1,1,1),(1,0,3)$ and $(0,1,m)$. One checks that summing the clockwise differences of values of $g$ going from a vertex to itself around any such color triple gives a zero contribution in $\mathbb{Z}_{3}$. Therefore, the value of $g[C](v)$ does not depend on the choice of the path from $A_{0}$ to $v$. That $(g[C]^{A},g[C]^{B},g[C]^{C})$ has the right boundary conditions is due to the fact that $C$ is regular. It remains to check the lozenge condition on $g[C]$ from Definition 2.1. Take any lozenge $l=(v^{1},v^{2},v^{3},v^{4})$ and suppose that $g[C](v^{2})=g[C](v^{4})$. Note that from Figure 3, the edge between $v^{2}$ and $v^{4}$ is always oriented from $v^{4}$ to $v^{2}$. The edge $e=(v^{4},v^{2})$ has color either $3$ or $m$. Since $C$ is a color map, the two faces adjacent to $e$ have either $(1,0,3)$ or $(0,1,m)$ colors. The face with vertices $(v^{1},v^{2},v^{4})$, respectively $(v^{3},v^{2},v^{4})$ is always direct, respectively reversed, see Figure 3. If $C(e)=3$, $g[C](v^{1})=g[C](v^{2})+1$ and $g[C](v^{3})=g[C](v^{2})+2$ whereas if $C(e)=m$, $g[C](v^{1})=g[C](v^{2})+2$ and $g[C](v^{3})=g[C](v^{2})+1$. In both cases, $\{g[C](v^{1}),g[C](v^{3})\}=\{g[C](v^{2})+1,g[C](v^{2})+2\}$ and thus $g[C]$ is a regular labeling.
Let us show that $g\mapsto C[g]$ maps a regular labeling $g$ to a color map. Since $g$ is regular, $C[g]$ also is by the same argument as above. Let us show that the only cyclic colors triples around any triangular face are $(0,0,0),(1,1,1),(1,0,3)$ and $(0,1,m)$. Take any triangular face and denote $X,Y,Z$ the clockwise differences of values of $g$. Then, $X+Y+Z=0[3]$ so that $(X,Y,Z)\in\{(1,1,1),(2,2,2),(0,1,2),(0,2,1)\}$ up to cyclic rotation. Note that we exclude $(0,0,0)$ by the lozenge condition in Definition 2.1 since no lozenge can have three vertices with equal values, for otherwise the second condition of Definition 2.1 is violated. These possible height differences give the clockwise colors $\{(1,1,1),(0,0,0),(3,1,0),(m,0,1)\}$ respectively. Therefore, $C[g]$ is a color map and by construction $g\mapsto C[g]$ is the inverse of $C\mapsto g[C]$. ∎
@@
The image $\Psi^{C}[(C,L)]$ of any limit hive $(C,L)\in H^{C}(\alpha,\beta,\gamma,\infty)$ can be extended by Remark 7.8 to a function $f:R_{d,n}\rightarrow\mathbb{R}$ such that by construction $\Psi^{C}[(C,L)]=f_{|R_{d,n}\setminus S^{C}}=f_{|Supp(g[C])}$. Let us check that $f\in P^{g[C]}_{\alpha,\beta,\gamma}$. By construction, the values of $f$ on $\partial R_{d,n}$ are as in Definition 2.4, see Figure 29. By definition of the extension, $f$ satisfies the equality constraints over any rigid lozenge in $R_{d,n}$. For any other lozenge $l=(v_{1},\dots,v_{4})$, the inequality $f(v_{2})+f(v_{4})\geq f(v_{1})+f(v_{3})$ is equivalent to the inequality (2c) of Definition 7.1.
@@
Figure 29. Boundary conditions on $P^{g[C]}_{\alpha,\beta,\gamma}$ induced by $\Psi^{C}$ on boundary conditions in Figure 27.
@@
For $C=C_{0}$, enumerate $e_{1},\ldots,e_{D}$ the horizontal edges in $C_{0}$ as in Lemma 6.24. Then, for $u\in\mathbb{R}^{D}$ and $v\in R_{d,n}$,
$[\Psi^{C_{0}}\circ\Phi^{C_{0}}(u)](v)=\sum_{e_{i}}(1-u_{i})+(d-v_{2})^{+}+\sum_{i=1}^{v_{2}}\beta_{i}$
@@
If $C$ is general, introduce for each hexagon rotation $C\rightarrow C^{\prime}$ given by an hexagon $h$ the map $\tilde{R}_{C\rightarrow C^{\prime}}:\tilde{P}^{C}_{\alpha,\beta,\gamma}\rightarrow\tilde{P}^{C^{\prime}}_{\alpha,\beta,\gamma}$ by
@@
    Let us describe how the hexagon rotation $C\rightarrow C^{\prime}$ maps $f$ to another function $f^{\prime}:R_{d,n}\rightarrow\mathbb{R}$. The value of the center vertex $c$ of $h$ is uniquely determined by the position of the rigid lozenge in $h$ and the values of $f$ on $\partial h$. Indeed, if $(v,v^{\prime},v^{\prime\prime})$ are the three other vertices of the rigid lozenge such that $C((c,v^{\prime\prime}))=m$, then $f(c)=f(v)+f(v^{\prime})-f(v^{\prime\prime})$. Note that $v,v^{\prime},v^{\prime\prime}\in\partial h$. For every vertex $u\in R_{d,n}$ other that the center vertex $c$ of $h$, we set $f^{\prime}(u)=f(u)$. In the hexagon rotation $C\rightarrow C^{\prime}$, the position of the the rigid lozenge changes and we set $f^{\prime}(c)=f(w)+f(w^{\prime})-f(w^{\prime\prime})$ where $w,w^{\prime},w^{\prime\prime}\in\partial h$ are the new vertices of the rigid lozenge which are located on $\partial h=\partial h$.
@@
    The map $\tilde{R}_{C\rightarrow C^{\prime}}(f)$ is defined as the restriction of $f^{\prime}$ of the previous step to $R_{d,n}\setminus S^{C^{\prime}}$.
Note that the map $\tilde{R}_{C\rightarrow C^{\prime}}$ is an affine bijection with integer coefficients whose inverse is given by $\tilde{R}_{C^{\prime}\rightarrow C}$. Let us check that the following diagram is commutative
@@
Let $H=(C,L)\in\tilde{H}^{C}(\alpha,\beta,\gamma,\infty)$ having an ABC hexagon $h$ with center vertex $c$. Denote by $C^{\prime}$ the color map obtained after any rotation $h\mapsto h^{\prime}$ and set $(C,L^{\prime})=Rot[C,C^{\prime}](C,L)$. For any vertex $u\neq c\in R_{d,n}$, one has $\tilde{R}_{C\rightarrow C^{\prime}}(\Psi^{C})(u)=\Psi^{C}(u)$. Moreover, if $u\neq c$ then one can find a path of edges from the south-east most vertex of $R_{d,n}$ to $u$ without any edge incident to $c$. Since the labels of these edges are not changed by $Rot[C\rightarrow C^{\prime}]$, $\Psi^{C^{\prime}}(L^{\prime})(u)=\Psi^{C}(L)(u)=\tilde{R}_{C\rightarrow C^{\prime}}(\Psi^{C}(L))(u)$ as desired. It remains to check that the same property holds for $u=c$. Denote $(v,v^{\prime},v^{\prime\prime})$, respectively $(w,w^{\prime},w^{\prime\prime})$ the vertices on $\partial h$ such that up to cyclic rotation $(v,c,v^{\prime},v^{\prime\prime})$, respectively $(w,c,w^{\prime},w^{\prime\prime})$, are vertices of the rigid lozenge in $h$, respectively $h^{\prime}$, and that $C((u,v^{\prime\prime}))=C^{\prime}((u,w^{\prime\prime}))=m$. Then,
@@
Since the values of $\Psi$ do not depend on the chosen path, let us choose the following four paths. Take any path $p=(e_{1},\dots,e_{r})\in(E_{n,d})^{r}$ from the south-east vertex $A_{0}$ to $w$ such that for each $1\leq i\leq r$, $e_{i}$ is not an interior edge of $h$. then,
@@
    To evaluate $\Psi^{C^{\prime}}(L^{\prime})(u)$, we append the edge $(w,u)$ to $p$,
@@
$\Psi^{C}(L)(w)$ $=\sum_{1\leq i\leq r}L(e_{i})+d$
$\Psi^{C}(L)(w^{\prime\prime})$ $=\sum_{1\leq i\leq r}L(e_{i})+L((w,w^{\prime\prime}))+d$
$\Psi^{C}(L)(w^{\prime})$ $=\sum_{1\leq i\leq r}L(e_{i})+L((w,w^{\prime\prime}))+L((w^{\prime\prime},w^{\prime}))+d$
$\Psi^{C^{\prime}}(L^{\prime})(u)$ $=\sum_{1\leq i\leq r}L^{\prime}(e_{i})+L^{\prime}((w,u))+d.$
@@
$L^{\prime}((w,u))=L((w,w^{\prime\prime}))+L((w^{\prime\prime},w^{\prime}))-L((w,w^{\prime\prime}))=L((w^{\prime\prime},w^{\prime})),$
@@
$L^{\prime}((w,u))=L((w^{\prime\prime},w^{\prime})).$
Notice that $(w,u),(w^{\prime\prime},w^{\prime})$ are two edges of the same type in the rigid lozenge in $C^{\prime}$ which implies that $L^{\prime}((w,u))=L^{\prime}((w^{\prime\prime},w^{\prime}))=L((w^{\prime\prime},w^{\prime}))$, where the last equality is due to the fact that $(w^{\prime},w^{\prime\prime})\in\partial h$ so that its label value is unchanged by $Rot[C,C^{\prime}]$. The commutativity of (7.23) is thus showed.
@@
By Theorem 3.8 we have
(7.24) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=\frac{sf(n-1)(2\pi)^{(n-1)(n-2)/2}\Delta(\mathrm{e}^{2i\pi\gamma})}{n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\lim_{N\rightarrow\infty}N^{-(n-1)(n-2)/2}c_{\lambda_{N},\mu_{N}}^{\nu_{N},d}.$
@@
(7.25) $\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]$ $=\frac{sf(n-1)(2\pi)^{(n-1)(n-2)/2}\Delta(\mathrm{e}^{2i\pi\gamma})}{n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{C}Vol[u\in\mathbb{R}^{D},\Phi^{C}[u]\in H^{C}(\alpha,\beta,\gamma,\infty)]$
(7.26) $=\frac{sf(n-1)(2\pi)^{(n-1)(n-2)/2}\Delta(\mathrm{e}^{2i\pi\gamma})}{n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{C}Vol(P^{g[C]}_{\alpha,\beta,\gamma})$
(7.27) $=\frac{sf(n-1)(2\pi)^{(n-1)(n-2)/2}\Delta(\mathrm{e}^{2i\pi\gamma})}{n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{g:R_{d,n}\rightarrow\mathbb{Z}_{3}\text{ regular}}Vol_{g}(P^{g}_{\alpha,\beta,\gamma}).$
@@
From the expression [36, Eq. (4.116)] proven in [15], we have